Integrated Nanosystems Research Facility


This weeks updates:
  • STS ICP is down due to a communication error. We’ve contacted STS and are waiting to hear back from them. The problem seems to be a bad module that will not communicate with the main PC.
  • Replaced the UPS system on the Cannon Stepper. System is now working within normal perimeters.
  • Finished repairing the CHA Evaporator Cryo Compressor and regenerated the system. Will do a dummy in the next day or so.
  • Allocated time to the FEI E-beam writing system. Reconfigured the computer cables so system is being checked out to see if the Nabady software will communicate with the stage.
  • The Gasonics asher is back online. The end of point error was corrected and is working within normal perimeters.
  • BMR Nitrite Oxide system, the gas manifold needs cleaned due to deposition being ununiformed and deposition rate has dropped.

Today will be the last day for the User Meetings in Calit rm 3355. Future meetings will be held at the same time in room EG 3311.

Jake will be out this week so there will be no user meeting this Wednesday. Please see cleanroom staff if you have any questions concerning the INRF or BiON.


A couple updates:

  • The STS system back on line and working fine.
  • We have not been able to get back to the CHA thermal evaporator.
  • EH&S has continued performing their annual hood inspection.
  • Gasonics system back online.
  • E-bean 1 evaporator went down due to operator error. Regened system Saturday, staff will check out system on Monday and should it be available.

INRF Staff

Just a reminder to please write your first & last name, date & time legibly in the log sheets. This is a small example of what i’m seeing and makes it very difficult to decipher. Thanks for your help!

Date - Time in - Time out